(2xy-y2+2x)dx-(2xy-x2-2y)dy=0


1
Expert's answer
2022-02-22T23:00:52-0500
"(2xy-y^2+2x)dx+(-2xy+x^2+2y)dy"

"M(x,y)=2xy-y^2+2x,"

"N(x,y)=-2xy+x^2+2y,"

"\dfrac{\partial M}{\partial y}=2x-2y=\dfrac{\partial N}{\partial x}"

We have the following system of differential equations to find the function "u(x, y)"

"\dfrac{\partial u}{\partial x}=2xy-y^2+2x"

"\dfrac{\partial u}{\partial y}=-2xy+x^2+2y"

By integrating the first equation with respect to "x," we obtain


"u(x, y)=\int (2xy-y^2+2x)dx+\varphi(y)"

"=x^2y-xy^2+x^2+\varphi(y)"

Substituting this expression for "u(x,y)" into the second equation gives us:


"x^2-2xy+\varphi'(y)=-2xy+x^2+2y"

"\varphi'(y)=2y"

By integrating the last equation, we find the unknown function "\varphi(y)"


"\varphi(y)=\int 2y dy=y^2-C"

So that the general solution of the exact differential equation is given by


"x^2y-xy^2+x^2+y^2=C"